OpenStudy (anonymous):

convert 16,629 to base 13

OpenStudy (lgbasallote):

\[13^x = 16,629\] \[\log_13 (16,629) = x\]

OpenStudy (lgbasallote):

\[\log_{13} (16,629) = x\]

jimthompson5910 (jim_thompson5910):

16629/13 = 1279 remainder 2 1279/13 = 98 remainder 5 98/13 = 7 remainder 7 7/13 = 0 remainder 7 Read the remainders in reverse order to get 7752 So 16629 base 10 = 7752 base 13